Sam and Rachel leave a park at the same time and ride home. Sam rides at 15 km/h and Rachel rides at 11 km/h. Sam gets home 12 minutes before Rachel. How long did it take Sam to ride home from the park?
Found 2 solutions by mananth, cher781:
Answer by mananth(16946)

Sam and Rachel leave a park at the same time and ride home. Sam rides at 15 km/h and Rachel rides at 11 km/h. Sam gets home 12 minutes before Rachel. How long did it take Sam to ride home from the park?
let Sam take t hours
Sam rate = 15 km/h
Rachel takes t+12 minutes = t+1/5 hours
Rachel rate = 11 km/h
Distance is the same
distance Rachel = distance Sam
d=rt
15t=11(t+(1/5))
15t=11((5t+1)/5)
multiply by 5
75t=11(5t+1)
75t=55t+11
20t=11
t=11/20 hours
11/20 * 60 = 33 minutes
Sam took 33 minutes








Answer by cher781(8)

Call Sam's time t. Then Rachel's time is t + 12.
d = st (distance = speed by time). They both covered the same distance, so you can write:
15t = 11(t + 12)
15t = 11t + 132
4t = 132
t = 33
It took Sam 33 minutes to get home.